# Using VLM Reasoning to Constrain Task and Motion Planning

### v1

@@ IV-2 Visual reasoning with VLMs
VIZ-COAST’s visual reasoning module (VRM) leverages the common-sense visual and spatial reasoning capabilities of VLMs to foresee potential downward refinement failures and infer necessary planning constraints in advance. As shown in Fig.˜2, this module takes four inputs, which condition each step of prompting: symbolic specifications of the problem and domain in PDDL, an image of the initial state of the scene, a text-based representation of the scene, and an example of a compliant output. Its output is a Python program which encodes the necessary task-level planning constraints to avoid downward refinement failures.
The VRM consists of a 4-step prompting procedure, as described in Fig.˜3. The first step is scene interpretation: from the inputs, generate a description of the scene in natural language. Intuitively, this forces the VLM to interpret the scene in a semantically meaningful way. In the second step, the VLM is asked to identify potential downward refinement issues and describe the necessary constraints to mitigate them in natural language. Then, the third prompt asks the VLM to encode these constraints formally as a Python program, using Z3’s API. The final step is proofreading, in which we simply ask the model to examine the output and ensure it is syntactically compliant and semantically consistent with the natural language description produced by the second prompt. A paraphrased example of a constraint script generated by the VRM which blocks items from being picked out of or placed into closed containers is shown in Fig.˜4.
@@ IV-3 Motion Planning
VIZ-COAST integrates with the COAST framework to perform downward refinement, which itself builds upon PDDLStream [4]’s use of streams for sampling continuous action parameters. As the purpose of our approach is to impose task-level constraints to prevent motion planning failures, we only make modifications at the task planning level. For motion planning, we adopt COAST’s stream-based approach directly without modification.
VIZ-COAST’s goal is to infer the necessary constraints zero-shot, before any planning failures occur. However, in the event that downward refinement still fails, we simply re-run the prompting procedure.
@@ V-A Domains
We evaluated VIZ-COAST on two TAMP domains, each involving a robot arm manipulating objects in a simulated PyBullet environment [37]:
  • Blocks: Rearrange a set of blocks from an initial configuration within a 3x3 grid such that the red block is in the center.
@@
Figure 5: The Block domain (left) requires the robot to rearrange blocks to move the red block to the center tile. The end-effector may only approach the grid from a single direction. The Containers domain (right) requires the robot to place the items in target containers, removing and replacing their lids as needed.
Each domain has a corresponding PDDL description for task planning. Importantly, this symbolic description neglects key considerations for downward refinement, which manifest as motion grounding failures. For each domain, we describe the failure modes and demonstrate that VIZ-COAST infers constraints that not only mitigate these failures more efficiently than prior approaches, but also generalize to the entire class of problem instances posed by the domain.
@@ V-B Baselines
VIZ-COAST enables faster planning than prior state-of-the-art through reduced replanning, as well as higher success rates. We demonstrate this by evaluating VIZ-COAST against 2 strong baselines as well as a naive VLM-based approach:
@@
Both COAST+FD and COAST+Z3 employ COAST’s approach of adding constraints upon encountering downward refinement failures. We use GPT-5 [38] as the VLM in VIZ-COAST’s VRM in experiments and as the VLM-only baseline, although any VLM can be used.
Figure 6: Planning times, refinement failures, and success rates for the Blocks domain, averaged over up to 5 trials. VIZ-COAST was prompted once on the first 1-block problem instance, producing constraints which generalized to all 35 unique problem instances containing up to 7 blocks. By preventing refinement failures and thereby avoiding replanning, VIZ-COAST achieves lower planning times than COAST+FD and COAST+Z3. The VLM-only baseline is far slower than all other approaches, and fails completely on large problem instances.
Figure 7: Planning times, refinement failures, and success rates for the Containers domain, averaged over up to 5 trials. VIZ-COAST was prompted once on the first 1-goal problem instance, and then reprompted on subsequent failures. VIZ-COAST is able to produce successful plans for all problem instances, whereas COAST+FD fails on all instances and COAST+Z3 only succeeds for some 1-goal instances. These failures arise because the task-level specification lacks geometric intuition—the precondition of removing the lid is not captured in the symbolic PDDL description.
V-C Block
The Block domain features a $3\times 3$ grid of tiles on which blocks are placed upright. In each problem instance, number of blocks and their initial configuration may vary, but the goal is always to place the single red block on the middle tile. When picking and placing blocks, the robot may only approach from a single direction; its end effector must be facing away from its base and towards the target block, as shown in Fig.˜5.
@@ V-C Block
Results for the Blocks domain are shown in Fig.˜6. Each method was evaluated on problem instances of up to 7 blocks, with 5 trials per number of blocks, for a total of 35 unique instances. Evaluation proceeded sequentially from 1-block to 7-block instances. Planning time is how long the task planner took to produce a successful high-level plan, and does not include motion planning time, which varied little between methods. Ref. Failures refers to the total number of times each method replanned due to a downward refinement failure. If for a particular problem instance, a method exceeded 40 refinement failures, the instance was deemed a failure. All numbers are averaged over 5 trials. To ensure fair comparison, only problem instances for which at least 3 out of 4 methods succeeded are included in the average. VIZ-COAST’s VRM constraint generation pipeline is called a single time on an example scene extracted from the first single-block problem instance. Constraints persist over trials until a motion grounding failure occurs, at which point the constraint generation pipeline is re-run.
Our experiments show that on the Blocks domain, VIZ-COAST produces successful constraints on the initial call, before any planning occurs, which generalize to all subsequent trials over all block counts. In other words, VIZ-COAST does not encounter any downward refinement failures on the Blocks domain. It is also worth noting that the problem instance from which the constraints were inferred only had one block, yet VIZ-COAST was still able to produce constraints that generalized to several blocks.
In contrast, the COAST-based baselines often encountered 5 or more refinement failures before finding a successful plan on instances with several blocks, resulting in longer total planning times. Interestingly, the VLM-only baseline performed very poorly, succeeding at less than half of all instances, with a 0% success rate on instances with 5 or more blocks.
@@ V-D Containers
The Containers domain consists of 3 containers, each with a movable lid, 6 movable items, and a plate. The robot is able to move the items and lids. In contrast with Blocks, Containers has a fixed initial state, instead varying the goal state between problem instances. Lids obstruct placing items into containers. This domain’s core challenge is that the PDDL domain specification lacks key geometric intuition: removing the lid is not explicitly expressed as a precondition for placing objects into containers. Constraints are required to ensure that objects can only be placed inside or picked out when the target container is open.
@@
Results for the Containers domain are shown in Fig.˜7. Each method was evaluated on problem instances of up to 6 goal predicates with 5 trials per number of goals for a total of 30 unique instances. Goal predicates are randomly generated, each assigning a specific item or lid to be in or on a specific container. VIZ-COAST’s constraint generation pipeline is initially called once before any planning occurs on an example scene extracted from the first problem instance. Upon encountering a downward refinement failure, a new example scene is extracted from the failed instance, and the VRM is called again. As in Blocks, the planning time is how long the task planner took to produce a high-level plan. Ref. Failures refers to the total number of times the method replanned due to a downward refinement failure. For this domain, we report the sum total number of refinement failures for each goal count. If for a particular problem instance a method failed to find a plan within one hour it was deemed a failure. All numbers are averaged over up to 5 trials. Note that COAST+FD failed to find a plan for any problem instance, and COAST+Z3 only succeeded on problem instances with one goal predicate. For goal counts with a 0% success rate, we cannot report planning times or refinement failures.
V-E Robustness
Method Success(%) Total ref. fails 5-block plan time(s)
VIZ-COAST 100 $\textbf{0.5}\pm\textbf{0.878}$ $0.328\pm 0.131$
Condensed 100 $2.0\pm 3.04$ $\textbf{0.296}\pm\textbf{0.090}$
COAST-FD $92.5$ $147.5\pm 104$ $1.257\pm 0.377$
COAST-Z3 8585 $126.5\pm 77.1$ $1.906\pm 0.404$
Table I: Comparison of success rates, average planning times, and average total downward refinement failure counts over 6 sequential 20-instance trials on the Block domain. VIZ-COAST’s 4-stage prompting procedure reliably generalizes across instances, averaging 4 times fewer refinement failures than Condensed (the single-prompt ablation of VIZ-COAST). In contrast, COAST-based baselines encounter frequent refinement failures, resulting in longer planning times.
We conduct an additional comparison in which we run 6 independent trials over 20 sequential problem instances in the Blocks domain for each method. To justify the VRM’s 4-stage prompting procedure, we introduce Condensed, an ablation of VIZ-COAST which produces constraints in just a single prompt from the same inputs.
Of the 6 trials, 3 consist of only 5-block instances, while the other 3 consist of 1 to 4-block instances. VIZ-COAST’s constraint generation pipeline is initially called once before any planning occurs on an example scene extracted from the first problem instance. Upon encountering a downward refinement failure, a new example scene is extracted from the failed instance, and the VRM is called again. Constraints persist between subsequent problem instances. We record each method’s success rate, average planning time on 5-block instances, and the average total number of failures encountered across all 20 instances in a trial. We note that this is not the number of refinement failures encountered on a single problem instance.
As shown in Tbl.˜I, VIZ-COAST’s constraints successfully generalize to all 20 problem instances on the first attempt, prior to any planning, on 4 out of 6 trials. On average, VIZ-COAST encounters $0.5$ motion grounding failures during each 20-instance trial before inferring generalizable constraints. In all trials for both VIZ-COAST and Condensed, any failures encountered were on the first problem instance.
Both VIZ-COAST and Condensed successfully produce downward-refinable task plans for all instances, taking $\sim$0.3 seconds for 5-block instances on average. However, Condensed encounters two motion grounding failures on average per trial, 4 times that of VIZ-COAST, indicating that our 4-stage prompting procedure enables more robust generalization.
Both COAST-based baselines are unable to solve some problem instances, and encounter significantly more downward refinement failures on average across trials. As a result, these methods take 3 to 4 times longer to solve 5-block instances on average, with VIZ-COAST providing a 74% speed-up over our fastest non-ablation baseline, COAST-FD.
@@ VI-A The Role of VLMs in TAMP
A key distinction between the Blocks and Containers domains is evidenced by the performance characteristics of the VLM-only baseline. In the Containers domain, the baseline successfully solved most problem instances across varying goal counts within a few iterations. This contrasts starkly with its performance in the Blocks domain, where it solved fewer than half of the problem instances, with success rates decreasing drastically as the number of blocks increased. We attribute this discrepancy to fundamental differences in the structure of the task abstraction of the two domains. Even with correct constraints, the Blocks domain requires long-horizon, non-ergodic search to find feasible plans, whereas in the Containers domain, once the key implications on action validity imposed by the container lids is captured (that is, objects cannot be placed in or picked from closed containers), successful plans tend to follow relatively simple patterns of pick-and-place. These results highlight both the strengths and limitations of VLMs in TAMP applications. The VLM-only baseline’s success in the Containers domain demonstrates that VLMs excel at leveraging common-sense visual reasoning to anticipate and avoid downward-refinement failures; conversely, their poor performance in the Blocks domain reveals their limitations in tasks that resolve search. By using VLMs exclusively for inference and imposition of constraints while delegating planning to a TAMP solver, VIZ-COAST exploits VLMs’ strengths while circumventing their inherent weaknesses.
VI-B The Advantage of Constraints in Task Planning
A primary factor contributing to the high failure rates of COAST-based baselines in the Containers domain is the requirement to introduce additional entities and predicates into the PDDL description to express constraints. This modification incurs significant overhead in search (due to the size of the grounding), as well as in the translation step to convert PDDL to SAS+ used by Fast Downward. In contrast, VIZ-COAST encodes constraints directly through the Z3 API, enabling the planner to utilize the original PDDL description while introducing minimal additional complexity. In the Containers domain, this overhead frequently determined whether a method could find a successful plan within seconds or fail entirely. This observation illustrates the computational advantage of SMT-based constraints while simultaneously highlighting a limitation of failure-driven constraint discovery: the process of encountering failures may itself be intractable.

### v2

@@ IV-2 Visual reasoning with VLMs
VIZ-COAST’s visual reasoning module (VRM) leverages the common-sense visual and spatial reasoning capabilities of VLMs to foresee potential downward refinement failures and infer necessary planning constraints in advance. This module takes five inputs, which condition each step of prompting: symbolic specification of the problem and domain in PDDL, an image of the initial state of the scene, a text-based geometric representation of the scene, a grasp specification, and an example of a compliant output. The grasp specification is a short natural-language description of how grasps and approach poses are sampled in the downstream streams-based grounding. Its output is a Python program which encodes the necessary task-level planning constraints to avoid downward refinement failures.
The VRM consists of an initial 2-step prompting procedure, followed by a feedback mechanism. For initial constraint generation, the VLM is first asked to identify potential downward refinement issues and articulate the necessary constraints to mitigate them in natural language. Then, the VLM is asked to encode these constraints as a Python program, using Z3’s API. A paraphrased example of a constraint script generated by the VRM is shown in Fig.˜3. In the event that planning still fails with these initial constraints, the VRM continues prompting the VLM in a multi-turn conversation to update the constraints using feedback from planning. There are three failure cases from planning: if the task planner fails to find a plan, the VLM is asked to make the constraints less restrictive; if the constraints raise errors during application, the VLM is given the error message and asked to debug the constraints; if a task plan is produced but results in downward refinement failure, the task plan and the stream that failed to be successfully sampled are written to a feedback cache, a temporary state-action block is added to the constraints to prevent revisiting the failed action, and planning is reattempted. If no plan succeeds after $K$ such attempts, the temporary blocks are removed and the VLM is asked to analyze the $K$ feedback entries and revise the constraints to account for these failures. This allows the VLM to reason over several unique failures of the current set of constraints to make a more informed correction.
@@ IV-3 Motion Planning
VIZ-COAST integrates with the COAST framework to perform downward refinement, which itself builds upon PDDLStream [garrett_pddlstream_2020]’s use of streams for sampling continuous action parameters. As the purpose of our approach is to impose task-level constraints to prevent motion planning failures, we only make modifications at the task planning level. For motion planning, we adopt COAST’s stream-based approach directly without modification.
@@ V-A Domains
We evaluated VIZ-COAST on three TAMP domains, each involving a robot arm manipulating objects in a simulated PyBullet environment [pybullet]:
  • Blocks: Rearrange blocks within a 3x3 grid such that the red block is in the center.
@@
  • Packing: Pack several items of varying shape into a container such that the lid may be placed on top.
Figure 4: The Blocks domain (left) requires the robot to rearrange blocks to move the red block to the center tile. The end-effector may only approach the grid from a single direction. The Containers domain (middle) requires the robot to place the items in target containers, removing and replacing their lids as needed. The Packing domain (right) requires the robot to cleverly pack all items into a container and close the lid.
Each domain has a corresponding PDDL description for task planning. Importantly, this symbolic description neglects key considerations for downward refinement, which manifest as motion grounding failures. For each domain, we describe the failure modes and demonstrate that VIZ-COAST infers constraints that not only mitigate these failures more efficiently than prior approaches, but also generalize across the class of problem instances posed by the domain.
@@ V-B Baselines
VIZ-COAST enables faster planning than prior state-of-the-art through reduced replanning, as well as higher success rates. We demonstrate this by evaluating VIZ-COAST against two strong baselines as well as two naive approaches.
@@
  • Incremental: A naive learning-from-failure method which adds a single state-action block to the SMT-based planner upon refinement failure, and iteratively extends task planning horizon, similar to IDTMP [dantam2018incremental].
@@
Both COAST+FD and COAST+Z3 employ COAST’s approach for adding constraints upon encountering downward refinement failures. We used Gemini 3 Flash on medium reasoning level as the VLM in VIZ-COAST’s VRM and the VLM-only baseline for baseline comparison experiments. A comparison of different VLM providers is presented in Section V-F.
Figure 5: Success rates, refinement failures, and planning times for the Blocks domain. VIZ-COAST was prompted once on the first 2-block problem, producing constraints which generalized to all 30 unique problems containing up to 7 blocks. VIZ-COAST achieves fewer refinement failures than COAST-based methods, while exhibiting comparable planning times.
Figure 6: Success rates, planning times, and refinement failures for the Containers domain. VIZ-COAST was prompted once on the first problem, producing constraints which generalized to all 30 unique problems containing up to 6 goal predicates. VIZ-COAST is able to produce successful plans for all problems, whereas COAST+FD fails on all problems and COAST+Z3 only succeeds for some 1-goal problems.
Figure 7: Success rates, refinement failures, and planning times for the Packing domain. VIZ-COAST was prompted once on the first 3-item problem to produce initial constraints, and then re-prompted on subsequent failures with feedback to revise constraints. VIZ-COAST achieves 100% success and fastest planning times on larger problem instances, whereas COAST-based methods become increasingly slower or fail completely. The failures of VIZ-COAST’s no-feedback ablation on 5-item problems shows the necessity of closing the constraint generation-planning loop in the VRM.
V-C Blocks
The Blocks domain features a $3\times 3$ grid of tiles on which blocks are placed upright. In each problem instance, the number of blocks and their initial configuration may vary, but the goal is always to place the single red block on the middle tile. When picking and placing blocks, the robot may only approach and grasp from a single direction; its end effector must be facing away from its base and towards the target block, as shown in Fig.˜4. This information is given to the VRM as part of the grasp specification.
@@ V-C Blocks
Results for the Blocks domain are shown in Fig.˜5. Each method was evaluated on problem instances of up to 2-7 blocks, with 5 problems per number of blocks, for a total of 30 unique instances. Evaluation proceeded sequentially from 2-block to 7-block instances. Planning time is how long the task planner took to produce a successful high-level plan, and does not include motion planning time, which varied little between methods. Ref. Failures refers to the total number of times each method replanned due to a downward refinement failure. If, for a particular problem, a method exceeded 100 refinement failures or a timeout of 600 seconds, the instance was deemed a failure. Metrics are averaged over up to all 5 problems at each number of goals. To ensure a fair comparison, only data points for which at least 3 methods succeeded were included in the average. VIZ-COAST’s VRM constraint generation pipeline is called a single time on an example scene extracted from the first single-block problem instance. Constraints persist over trials until a motion grounding failure occurs, at which point the VRM’s failure feedback mechanism updates the constraints.
Our experiments show that on the Blocks domain, VIZ-COAST produces successful constraints on the initial call, before any planning occurs, which generalize to all subsequent trials over all block counts. VIZ-COAST’s feedback mechanism used re-prompt threshold $K=1$, such that any failure would immediately trigger a constraint update. In other words, VIZ-COAST does not encounter any downward refinement failures on the Blocks domain. In contrast, COAST-based baselines often encountered as many as 10 failures before finding a successful plan on many-block instances. Though COAST-based methods are faster on simpler problems, their planning times show an exponential upward trend as the number of blocks increases.
@@ V-D Containers
The Containers domain consists of 3 containers with removable lids, 6 movable items, and a static plate where lids may be placed. In contrast with Blocks, Containers has a fixed initial state, instead varying the goal state between problem instances. Lids obstruct placing items into containers. Objects are grasped and approached from above. This domain’s core challenge is that the PDDL domain specification lacks key geometric intuition: removing the lid is not explicitly expressed as a precondition for placing objects into containers. Constraints are required to ensure that objects can only be placed inside or picked out when the target container is open.
@@
Results for the Containers domain are shown in Fig.˜6. Each method was evaluated on a fixed set of problem instances of 1-6 goal predicates with 5 problems per number of goals for a total of 30 unique instances. Each problem’s goal is a randomly generated set of predicates, each assigning a specific item or lid to be in or on a specific container. VIZ-COAST was able to find generalizable constraints on the first attempt. Although it encountered a small number of refinement failures on subsequent problems, successful plans were found within a constant number ($K=5$) incremental iterations. COAST+FD failed to find a plan for any problem instance, and COAST+Z3 only succeeded on problem instances with one goal predicate. For goal counts with a 0% success rate, we cannot report planning times or failure count.
V-E Packing
The Packing domain consists of 1 container with a lid and 3 kinds of packable items: slabs, big blocks, and small blocks. In each problem, all items start on the table, and the goal is to pack them all into the container and then place the lid on top. The symbolic domain description does not account for container’s limited internal volume, thus without constraints a trivial task plan is to pack all objects directly on the container floor. However, this is often geometrically infeasible as 1 slab occupies the entire floor. Constraints must be added to force the task planner to stack items in a configuration such that they will all fit horizontally without exceeding the vertical limit, so that the lid can be placed on top.
Results for the Packing domain are shown in Fig.˜7. Each method was evaluated on a fixed set of problem instances of 3, 4, 5, and 6 packable items where problem counts were 7, 8, 8, and 7, respectively, for a total of 30 unique instances. VIZ-COAST’s feedback mechanism used re-prompt threshold $K=5$.
Our results show that VIZ-COAST consistently achieves the lowest planning times on problem instances with 5 or 6 items to pack. In contrast, COAST+FD is fast on easier problems, but quickly begins to slow down and fail as complexity increases, failing completely on 6-item problems. COAST+Z3 fails on all problems. Remarkably, the VLM-only baseline achieves 100% success and a comparable number of refinement failures to VIZ-COAST, although it takes much longer to plan.
Unlike in Blocks and Containers, VIZ-COAST was unable to find fully-generalizable constraints on the initial attempt, requiring continued prompting using feedback from subsequent problems to correct constraints. Thus, we compare against an ablation of VIZ-COAST which does not use the feedback mechanism, instead starting a new conversation from scratch upon refinement failure. This ablation failed on several 5-item problems, and encountered more refinement failures than VIZ-COAST, demonstrating the benefit of using feedback from planning in the VRM.
V-F VRM Robustness
VLM Success Prompting(s) Ref. Failures
Gemini3-Pro-High 100% $94.09\pm 17.24$ ${\bf 0.00}\pm{\bf 0.00}$
Gemini3-Pro-Med 100% $54.79\pm 34.60$ $0.05\pm 0.22$
Gemini3-Flash-Med 100% $61.06\pm 11.24$ $0.15\pm 0.65$
Gemini3-Flash-Low 100% ${\bf 14.94}\pm{\bf 3.33}$ $0.20\pm 0.40$
GPT5.2-High 100% $113.68\pm 48.56$ $0.35\pm 0.96$
GPT5.2-Med 100% $53.91\pm 13.07$ $0.60\pm 1.36$
GPT5.2-Low 100% $29.25\pm 6.09$ $0.70\pm 0.95$
Qwen3.5-Plus 100% $184.36\pm 129.12$ $1.05\pm 2.58$
Qwen3.5-Flash $85\%$ $56.93\pm 49.31$ $4.59\pm 5.26$
Table I: Comparison of success rates, average time-per-prompt, and average total downward refinement failure counts of various VLM models over 20 independent trials of 4 sequential 4-6 block problems in the Block domain. Gemini models performed best, with refinement failures decreasing as reasoning level increased.
To evaluate the consistency with which the VRM infers constraints that successfully prevent downward refinement failure, we conduct an additional comparison in which we run 20 independent trials over 4 sequential problem instances in the Blocks domain for various VLM configurations.
We evaluated 5 VLM models at various reasoning levels across 3 providers. We report each VLM’s success rate (Success), average time per constraint update (Prompting), and the average total number of refinement failures encountered over all problems in a trial (Ref. Failures). Metrics are averaged over runs in which all problems succeeded. We set the VRM’s re-prompting threshold $K=1$ so that refinement failures would immediately trigger a constraint update.
The Gemini family of models performed best overall, with Gemini-3-Flash-Low yielding fewer refinement failures on average than GPT-5.2-High, despite taking nearly one-tenth of the time on average. Increasing chain-of-thought reasoning corresponded with a reduction in refinement failures, indicating that VLMs’ spatial reasoning capabilities improve with reasoning budget. All models achieved a 100% success rate except for Qwen-3.5-Flash.
@@ VI-A The Role of VLMs in TAMP
The motivating principle behind TAMP is that efficient planning requires the right abstraction. We posit that the success of VIZ-COAST stems from assigning VLMs a role aligned with their strengths: synthesizing constraints to augment the symbolic abstractions used in TAMP. The disparity in performance of the VLM-only baseline across domains illustrates this point. In Containers and Packing, this baseline achieved a 100% success rate, whereas in Blocks its success rate dropped as the number of blocks increased. We attribute this discrepancy to differences in what makes each domain challenging: in Containers and Packing, once key geometric constraints are recognized, planning becomes straightforward, whereas the Blocks domain still requires long-horizon, non-ergodic search. VLMs demonstrate strong geometric intuition but remain inefficient at this precise search. By isolating the VLM to the task of inferring constraints while delegating search to a symbolic planner, VIZ-COAST leverages the geometric intuition of VLMs while retaining the efficiency and precision of classical planning.
More broadly, VIZ-COAST belongs to a growing class of approaches that leverage VLMs to make TAMP more tractable. These approaches suggest that learned commonsense models can complement classical planning systems by leveraging knowledge that is difficult to encode symbolically. In future work, it would be interesting to explore integrating constraint synthesis with complementary strategies such as VLM-TAMP, which uses VLMs to propose intermediate subgoals which reduce the search space for long-horizon tasks.
VI-B The Strengths and Limits of VLMs in Constraint Synthesis
Table I shows that VIZ-COAST performs robustly across a variety of VLM providers, models, and reasoning budgets. In general, increasing the chain-of-thought reasoning budget leads to more consistent inference of correct constraints, suggesting that deeper reasoning improves the model’s ability to extract the geometric structure required for downward refinability. At the same time, fast models such as Gemini-3-Flash still achieve strong performance while requiring as little as 15 seconds to generate constraints.
VIZ-COAST aims first to act as a preventative method, proactively foreseeing and avoiding refinement failures before planning. However, as demonstrated in the Packing domain, VLMs are not always able to infer all necessary constraints a priori. Thus, in more challenging domains where initial constraints fail, VIZ-COAST leverages feedback from planning to update constraints. The Packing domain shows that this feedback is not merely helpful for recovering from occasional errors, but can be necessary for arriving at the correct insight. In particular, the no-feedback ablation was unable to solve several problem instances despite repeated attempts, whereas the feedback-enabled version succeeded. The ablation also encountered more refinement failures, suggesting that VIZ-COAST’s ability to reason over multiple problems produces more generalizable constraints.
VI-C The Advantage of SMT-based Constraints
A primary factor contributing to the high failure rates of COAST-based baselines in the Containers and Packing domains is the mechanism used to encode constraints. In COAST, expressing constraints requires introducing additional entities and predicates into the PDDL domain description, including timesteps and auxiliary logging and blocking predicates. These substantially increase the complexity of the domain representation. In other words, COAST’s mechanism for encoding constraints enables failure-driven constraint discovery at the expense of the utility of its abstraction.
In contrast, VIZ-COAST encodes constraints directly through the Z3 API as additional clauses in the intermediate SMT representation. This allows the planner to operate on the original PDDL description while incorporating constraints with minimal additional complexity. In the Containers and Packing domains, this difference in representation often determined whether a planner could find a solution within seconds or fail entirely.